# On relative tractor bundles

### v1

@@ 3.2. The relation to the Lie bracket
(1) We first observe that by Theorem 7, the condition on $\tau$ in (1) is equivalent to the fact that $[\xi,\eta]\in\Gamma(T^{i^{\prime}}_{P}M)$ for any $\xi\in\Gamma(T_{\rho}M)$ and $\eta\in\Gamma(T^{i^{\prime}}_{P}M)$. To prove necessity of this condition, we assume that $T^{i^{\prime}}N\subset TN$ is a subbundle with the required property. Then for a local section $\underline{\eta}$ of $T^{i^{\prime}}N$ we can first pull back to a local section of $T^{i^{\prime}}_{P}M/T_{\rho}M$ and then choose a representative section $\eta\in\Gamma(T^{i^{\prime}}_{P}M)$ for this. This means that $T_{y}\psi(\eta(Y))=\underline{\eta}(\psi(y))$ on the domain of definition of $\eta$, so the vector fields $\eta$ and $\underline{\eta}$ are $\psi$-related. Moreover, $\xi\in\Gamma(T_{\rho}M)$ is of course $\psi$-related to the zero vector field on $N$. This shows that $[\xi,\eta]$ has to be $\psi$-related to the zero vector field, too, and hence $[\xi,\eta]\in\Gamma(T_{\rho}M)$. Starting with a local frame of $T^{i^{\prime}}N$ and adding a local frame of $T_{\rho}M$ we conclude that there are local frames for $T^{i^{\prime}}_{P}M$ consisting of vector fields $\eta$ such that $[\xi,\eta]\in\Gamma(T_{\rho}M)$ for any $\xi\in\Gamma(T_{\rho}M)$, which clearly implies that the bracket with $\xi$ sends any section of $T^{i^{\prime}}_{P}M$ to a section of this subbundle.
@@
To prove sufficiency of the condition, put $n=\dim(M)$, $\ell=\text{rank}(T^{i^{\prime}}_{P}M)$ and $k=\text{rank}(T_{\rho}M)$. We adapt an argument from the proof of the Frobenius theorem in [Michor:topics, Section 3.18]. Recall that given a point $x\in U$ we can find a Frobenius chart $(V,v)$ (for the involutive distribution $T_{\rho}M$) around $x$ such that $V\subset U$. This means that $v(V)\subset\mathbb{R}^{n}$ can be written as $W_{1}\times W_{2}$ for connected open subsets in $\mathbb{R}^{k}$ and $\mathbb{R}^{n-k}$, such that for each $a\in W_{2}$, the pre-image $v^{-1}(W_{1}\times\{a\})$ is an integral submanifold for the distribution $T_{\rho}M$. Since $\psi$ is a submersion, $\psi(V)\subset N$ is open and by construction the $\mathbb{R}^{n-k}$-component of $v$ descends to a map $\psi(V)\to W_{2}$. This is smooth by the universal property of surjective submersions and a local diffeomorphism by construction. Shrinking $V$ if necessary, this becomes a diffeomorphism and hence can be used as a chart map on $N$. We implement this by viewing the local coordinates $v^{i}$ for $i=k+1,\dots,n$ to be defined both on $V$ and on $\psi(V)$, and then in these coordinates $\psi$ is given by $(v^{1},\dots,v^{n})\mapsto(v^{k+1},\dots,v^{n})$.
@@
Let us denote by $\partial_{i}$ the coordinate vector field $\frac{\partial}{\partial v^{i}}$ for each $i$. Then by construction $\partial_{1},\dots,\partial_{k}$ form a local frame for $T_{\rho}M$ and, shrinking $V$ if necessary, we can extend them by $\eta_{k+1},\dots,\eta_{\ell}\in\Gamma(T^{i^{\prime}}_{P}M|_{V})$ to a local frame for $T^{i^{\prime}}_{P}M|_{V}$. Expanding these sections in terms of the $\partial_{i}$, we can leave out all summands with $i\leq k$, so we can assume that there are smooth functions $a_{ij}:V\to\mathbb{R}$ for $1\leq i\leq n-k$ and $1\leq j\leq\ell-k$ such that $\eta_{j+k}=\sum_{i=1}^{n-k}a_{ij}\partial_{k+i}$. Viewing $(a_{ij})$ as a matrix of size $(n-k)\times(\ell-k)$, it follows from linear independence of the fields $\eta_{j}$ that the matrix $(a_{ij}(y))$ has rank $\ell-k$ for each $y\in V$ and hence has $\ell-k$ linearly independent rows. Specializing to $y=x$ we can assume (permuting coordinates in $W_{2}$ if necessary) that the first $\ell-k$ rows of $(a_{ij}(x))$ are linearly independent and hence the top $(\ell-k)\times(\ell-k)$ sub-matrix is invertible. Again shrinking $V$ if necessary, we can assume that the corresponding sub-matrix in $(a_{ij}(y))$ is invertible for any $y\in V$. Hence we find smooth functions $b_{ij}:V\to\mathbb{R}$ for $1\leq i,j\leq\ell-k$ such that $\sum_{r}a_{ir}(y)b_{rj}(y)=\delta_{ij}$ for any $i,j$ and any $y\in V$.
@@
Now we define $\tilde{\eta}_{k+j}:=\sum_{i=1}^{\ell-k}b_{ij}\eta_{k+i}$ for $i=1,\dots,\ell-k$. By construction, these a smooth sections of $T^{i^{\prime}}_{P}M$ which in each point span the same subspace as the $\eta$’s and hence also extend $\partial_{1},\dots,\partial_{k}$ to a local frame of $T^{i^{\prime}}_{P}M$ on $V$. Plugging in the expansions of the $\eta$’s in terms of the coordinate vector fields, one immediately concludes that $\tilde{\eta}_{k+j}=\partial_{k+j}+\sum_{i=\ell+1}^{n}c_{ij}\partial_{i}$ for some smooth functions $c_{ij}:V\to\mathbb{R}$. Assuming that $[\xi,\eta]\in\Gamma(T^{i^{\prime}}_{P}M)$ for any $\xi\in\Gamma(T_{\rho}M)$ and $\eta\in\Gamma(T^{i^{\prime}}_{P}M)$ we conclude that for each $i=1,\dots,k$, $[\partial_{i},\tilde{\eta}_{k+j}]$ can be expanded in as a smooth linear combination of $\partial_{1},\dots,\partial_{k},\tilde{\eta}_{k+1},\dots,\tilde{\eta}_{\ell}$. But this bracket is given by $\sum_{r\geq\ell+1}(\partial_{i}\cdot c_{rj})\partial_{r}$ and in view of the form of the $\tilde{\eta}$’s this is only possible if $[\partial_{i},\tilde{\eta}_{k+j}]=0$ for any $i=1,\dots k$ and $j=1,\dots,\ell-k$. But this implies that all the functions $c_{rj}$ satisfy $\partial_{i}\cdot c_{rj}=0$ for all $i$, whence they are independent of the coordinates $v^{1},\dots,v^{k}$ and hence constant along the fibers of $\psi$.
@@
(2) The first statement is obvious by part (1). Under the additional assumption on $\tau$, part (2) of Theorem 7 shows that for a representative $\eta\in\Gamma(T^{i^{\prime}}_{P}M)$ for $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$ one obtains $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$ as the projection of $[\xi,\eta]$. Hence what we have to prove here is that $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}\sigma=0$ if and only if for one or equivalently any representative $\eta$ for $\nabla$ we get $[\xi,\eta]\in\Gamma(T^{i^{\prime}+1}_{P}M)$ for any $\xi\in\Gamma(T_{\rho}M)$. Now we can run the argument with a Frobenius chart as in the proof of (1) in two steps. First, we extend the local frame $(\partial_{i})_{i=1,\dots,k}$ for $T_{\rho}M$ by vector fields $\eta_{k+1},\dots,\eta_{k+\ell^{\prime}}$ to a local frame of $T^{i^{\prime}+1}_{P}M$ in such a way that $\eta_{k+j}=\partial_{k+j}+\sum_{r>k+\ell^{\prime}}a_{r}\partial_{r}$ for some smooth functions $a_{r}$. Then we can further extend by $\tilde{\eta}_{k+\ell^{\prime}+1},\dots,\tilde{\eta}_{k+\ell}$ to a local frame for $T^{i^{\prime}}_{P}M$. As before, we can obviously assume that the $\tilde{\eta}$’s are smooth linear combinations of the coordinate vector fields $\partial_{r}$ with $r>k+\ell^{\prime}$ only. As in part (1), we can then modify these to fields $\eta_{k+\ell^{\prime}+1},\dots,\eta_{k+\ell}$ such that $\eta_{k+\ell^{\prime}+j}=\partial_{k+\ell^{\prime}+j}+\sum_{r>k+\ell}c_{rj}\partial_{r}$ for smooth functions $c_{rj}$.
@@
In particular, as in part (1) the vector fields $\eta_{k+\ell^{\prime}+1},\dots,\eta_{k+\ell}$ project to a local frame for $\mathcal{V}_{i^{\prime}}M$ that descends to a local frame for $T^{i^{\prime}}N/T^{i^{\prime}+1}N$. Now in the domain of our Frobenius chart, take a section of $\mathcal{V}_{i^{\prime}}M$ and a representative in $\Gamma(T^{i^{\prime}}_{P}M)$. Expanding this in terms of our local frame, we can drop all summands except the last $\ell-\ell^{\prime}$ ones without changing the projection to $\Gamma(\mathcal{V}_{i^{\prime}}M)$. Hence we find a representative of the form $\eta=\sum_{j=1}^{\ell-\ell^{\prime}}c_{j}\eta_{k+\ell^{\prime}+j}$ and $[\partial_{i},\eta]=\sum_{j=1}^{\ell-\ell^{\prime}}(\partial_{i}\cdot c_{j})\eta_{k+\ell+j}$. The form of $\eta_{k+\ell+j}$ readily implies that this is a section of $T^{i^{\prime}+1}M$ if and only if it is zero and hence to $\partial_{i}\cdot c_{j}=0$ for any $i$ and $j$. But this is equivalent to each $c_{j}$ being independent of the first $k$ coordinates. Hence $\eta$ descends to a local section of $T^{i^{\prime}}N/T^{i^{\prime}+1}N$ whose pullback visibly coincides with the original section of $\mathcal{V}_{i^{\prime}}M$. Conversely, in our frame, any such pullback has coordinate functions that descend to $N$. ∎
3.3. The distinguished class of relative tractor bundles
@@
In the setting of §2.3 let us assume that we have given a parabolic geometry $(p:\mathcal{G}\to M,\omega)$ whose torsion $\tau$ satisfies $\tau(T_{\rho}M,T^{i^{\prime}}_{P}M)\subset T^{i^{\prime}}_{P}M$ for all $i^{\prime}\leq 0$. Then $T_{\rho}M$ is involutive and by Theorem 8 for any local leaf space $\psi:U\to N$ for the induced foliation with connected fibers, we get a canonical filtration of the tangent bundle by the smooth subbundles $T^{i^{\prime}}N$. As usual, one can form the associated graded of this filtration, which has the form
@@
$\operatorname{gr}(TN)=\oplus_{i^{\prime}}\operatorname{gr}_{i^{\prime}}(TN)\text{\ with\ }\operatorname{gr}_{i^{\prime}}(TN)=T^{i^{\prime}}N/T^{i^{\prime}+1}N\text{\ and\ }T^{0}N=N\times\{0\}.$
@@
Performing natural constructions with vector bundles, one always obtains induced filtrations of the resulting bundles, and forming the associated graded is compatible with the constructions. For example, the natural filtration of the cotangent bundle $T^{*}N$ has positive indices with the $j$th filtration component being the annihilator of $T^{-j+1}N$ so $TN=T^{1}N\supset T^{2}N\supset\dots$. For the associated graded, this readily implies that $\operatorname{gr}_{j}(T^{*}N)$ is dual to $\operatorname{gr}_{-j}(TN)$. For filtered bundles $(E,E^{i})$ and $(F,F^{j})$ the components of the filtration on the tensor product $E\otimes F$ are defined by letting $(E\otimes F)^{\ell}$ be the span of the component $E^{i}\otimes F^{j}$ with $i+j=\ell$. For the associated graded, this means that $\operatorname{gr}_{\ell}(E\otimes F)=\oplus_{i+j=\ell}\operatorname{gr}_{i}(E)\otimes\operatorname{gr}_{j}(F)$. See Section 3.1.1 of [book] for more details.
@@
Corollary 9 (to Theorems 7 and 8).
In the setting of §2.3 consider a parabolic geometry $(p:\mathcal{G}\to M,\omega)$ whose torsion $\tau$ satisfies $\tau(T_{\rho}M,T^{i^{\prime}}_{P}M)\subset T^{i^{\prime}}_{P}M$ for all $i^{\prime}\leq 0$. Let $\psi:U\to N$ be a local leaf space for the involutive distribution $T_{\rho}M=T^{0}_{P}M$ with connected fibers.
(1) For any $i^{\prime}$ we get $\mathcal{V}_{i^{\prime}}M|_{U}=\psi^{*}(\operatorname{gr}_{i^{\prime}}(TN))$. Applying any tensorial construction to the bundles $\mathcal{V}_{i^{\prime}}M$, one obtains a relative tractor bundle whose restriction to $U$ can be naturally identified with the pullback of the the same construction applied to the bundles $\operatorname{gr}_{i^{\prime}}(TN)$.
(2) Suppose that the geometry even satisfies $\tau(T_{\rho}M,T^{i^{\prime}}_{P}M)\subset T^{i^{\prime}+1}_{P}M$ for all $i^{\prime}<0$. Then for any relative tractor bundle constructed as in (1) a local section over $U$ is parallel for the relative tractor connection if and only if it is a pullback of a section of the corresponding bundle over $N$.
@@
(1) The first part immediately follows from Theorems 7 and 8. Applying a tensorial construction vector bundles associated to a principal bundle, one always obtains the bundle associated to the representation obtained from the corresponding construction on the inducing representations. Hence any tensorial construction with the bundles $\mathcal{V}_{i^{\prime}}M$ leads to a relative tractor bundle. The last part then follows from the compatibility of tensorial constructions with the passage from filtered vector bundles to associated graded bundles as discussed above and with pullbacks.
@@
(2) Recall the construction of the relative tractor connection in terms of the fundamental derivative (or its relative analog) and the infinitesimal representation as described in §3.1. Naturality of the fundamental derivative (see Section 1.5.8 of [book]) then implies that applying a tensorial construction to the bundles $\mathcal{V}_{i^{\prime}}M$, the relative tractor connection on the resulting relative tractor bundle coincides with the connection induced by the relative tractor connection on the bundles $\mathcal{V}_{i^{\prime}}M$. Knowing this, the general claim follows from Theorem 8. ∎
3.4. Example: Legendrean contact structures

### v2

@@ 2. $(2)$
(1) We first observe that by Theorem 3.3, the condition on $\tau$ in (1) is equivalent to the fact that $[\xi,\eta]\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ for any $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$ and $\eta\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$. To prove necessity of this condition, we assume that $T^{i^{\prime}}N\subset TN$ is a subbundle with the required property. Then for a local section $\underline{\eta}$ of $T^{i^{\prime}}N$ we can first pull back to a local section of $T^{i^{\prime}}_{P}M/T_{\rho}M$ and then choose a representative section $\eta\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ for this. This means that $T_{y}\psi(\eta(y))=\underline{\eta}(\psi(y))$ on the domain of definition of $\eta$, so the vector fields $\eta$ and $\underline{\eta}$ are $\psi$-related. Moreover, $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$ is of course $\psi$-related to the zero vector field on $N$. This shows that $[\xi,\eta]$ has to be $\psi$-related to the zero vector field, too, and hence $[\xi,\eta]\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$. Starting with a local frame of $T^{i^{\prime}}N$ and adding a local frame of $T_{\rho}M$, we conclude that there are local frames for $T^{i^{\prime}}_{P}M$ consisting of vector fields $\eta$ such that $[\xi,\eta]\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$ for any $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$, which clearly implies that the bracket with $\xi$ sends any section of $T^{i^{\prime}}_{P}M$ to a section of this subbundle.
@@
To prove sufficiency of the condition, put $n=\dim(M)$, $\ell=\text{rank}\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ and $k=\text{rank}\bigl{(}T_{\rho}M\bigr{)}$. We adapt an argument from the proof of the Frobenius theorem in [19, Section 3.18]. Recall that given a point $x\in U$ we can find a Frobenius chart $(V,v)$ (for the involutive distribution $T_{\rho}M$) around $x$ such that $V\subset U$. This means that $v(V)\subset\mathbb{R}^{n}$ can be written as $W_{1}\times W_{2}$ for connected open subsets in $\mathbb{R}^{k}$ and $\mathbb{R}^{n-k}$, such that for each $a\in W_{2}$, the pre-image $v^{-1}(W_{1}\times\{a\})$ is an integral submanifold for the distribution $T_{\rho}M$. Since $\psi$ is a submersion, $\psi(V)\subset N$ is open and by construction the $\mathbb{R}^{n-k}$-component of $v$ descends to a map $\psi(V)\to W_{2}$. This is smooth by the universal property of surjective submersions and a local diffeomorphism by construction. Shrinking $V$ if necessary, this becomes a diffeomorphism and hence can be used as a chart map on $N$. We implement this by viewing the local coordinates $v^{i}$ for $i=k+1,\dots,n$ to be defined both on $V$ and on $\psi(V)$, and then in these coordinates $\psi$ is given by $\bigl{(}v^{1},\dots,v^{n}\bigr{)}\mapsto\bigl{(}v^{k+1},\dots,v^{n}\bigr{)}$.
@@
Let us denote by $\partial_{i}$ the coordinate vector field $\frac{\partial}{\partial v^{i}}$ for each $i$. Then by construction $\partial_{1},\dots,\partial_{k}$ form a local frame for $T_{\rho}M$ and, shrinking $V$ if necessary, we can extend them by $\eta_{k+1},\dots,\eta_{\ell}\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M|_{V}\bigr{)}$ to a local frame for $T^{i^{\prime}}_{P}M|_{V}$. Expanding these sections in terms of the $\partial_{i}$, we can leave out all summands with $i\leq k$, so we can assume that there are smooth functions $a_{ij}\colon V\to\mathbb{R}$ for $1\leq i\leq n-k$ and $1\leq j\leq\ell-k$ such that $\eta_{j+k}=\sum_{i=1}^{n-k}a_{ij}\partial_{k+i}$. Viewing $(a_{ij})$ as a matrix of size $(n-k)\times(\ell-k)$, it follows from linear independence of the fields $\eta_{j}$ that the matrix $(a_{ij}(y))$ has rank $\ell-k$ for each $y\in V$ and hence has $\ell-k$ linearly independent rows. Specializing to $y=x$, we can assume (permuting coordinates in $W_{2}$ if necessary) that the first $\ell-k$ rows of $(a_{ij}(x))$ are linearly independent and hence the top $(\ell-k)\times(\ell-k)$ submatrix is invertible. Again shrinking $V$ if necessary, we can assume that the corresponding submatrix in $(a_{ij}(y))$ is invertible for any $y\in V$. Hence, we find smooth functions $b_{ij}\colon V\to\mathbb{R}$ for $1\leq i,j\leq\ell-k$ such that $\sum_{r}a_{ir}(y)b_{rj}(y)=\delta_{ij}$ for any $i$, $j$ and any $y\in V$.
@@
Now we define $\tilde{\eta}_{k+j}:=\sum_{i=1}^{\ell-k}b_{ij}\eta_{k+i}$ for $i=1,\dots,\ell-k$. By construction, these are smooth sections of $T^{i^{\prime}}_{P}M$ which in each point span the same subspace as the $\eta$’s and hence also extend $\partial_{1},\dots,\partial_{k}$ to a local frame of $T^{i^{\prime}}_{P}M$ on $V$. Plugging in the expansions of the $\eta$’s in terms of the coordinate vector fields, one immediately concludes that $\tilde{\eta}_{k+j}=\partial_{k+j}+\sum_{i=\ell+1}^{n}c_{ij}\partial_{i}$ for some smooth functions $c_{ij}\colon V\to\mathbb{R}$. Assuming that $[\xi,\eta]\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ for any $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$ and $\eta\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ we conclude that for each $i=1,\dots,k$, $[\partial_{i},\tilde{\eta}_{k+j}]$ can be expanded in as a smooth linear combination of $\partial_{1},\dots,\partial_{k},\tilde{\eta}_{k+1},\dots,\tilde{\eta}_{\ell}$. But this bracket is given by $\sum_{r\geq\ell+1}(\partial_{i}\cdot c_{rj})\partial_{r}$ and in view of the form of the $\tilde{\eta}$’s this is only possible if $[\partial_{i},\tilde{\eta}_{k+j}]=0$ for any $i=1,\dots k$ and $j=1,\dots,\ell-k$. But this implies that all the functions $c_{rj}$ satisfy $\partial_{i}\cdot c_{rj}=0$ for all $i$, whence they are independent of the coordinates $v^{1},\dots,v^{k}$ and hence constant along the fibers of $\psi$.
@@
(2) The first statement is obvious by part (1). Under the additional assumption on $\tau$, part (2) of Theorem 3.3 shows that for a representative $\eta\in\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$ for $\sigma\in\Gamma(\mathcal{V}_{i^{\prime}}M)$ one obtains $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}_{\xi}\sigma$ as the projection of $[\xi,\eta]$. Hence, what we have to prove here is that $\nabla^{\rho,\mathcal{V}_{i^{\prime}}}\sigma=0$ if and only if for one or equivalently any representative $\eta$ for $\nabla$ we get $[\xi,\eta]\in\Gamma\bigl{(}T^{i^{\prime}+1}_{P}M\bigr{)}$ for any $\xi\in\Gamma\bigl{(}T_{\rho}M\bigr{)}$. Now we can run the argument with a Frobenius chart as in the proof of (1) in two steps. First, we extend the local frame $(\partial_{i})_{i=1,\dots,k}$ for $T_{\rho}M$ by vector fields $\eta_{k+1},\dots,\eta_{k+\ell^{\prime}}$ to a local frame of $T^{i^{\prime}+1}_{P}M$ in such a way that $\eta_{k+j}=\partial_{k+j}+\sum_{r>k+\ell^{\prime}}a_{r}\partial_{r}$ for some smooth functions $a_{r}$. Then we can further extend by $\tilde{\eta}_{k+\ell^{\prime}+1},\dots,\tilde{\eta}_{k+\ell}$ to a local frame for $T^{i^{\prime}}_{P}M$. As before, we can obviously assume that the $\tilde{\eta}$’s are smooth linear combinations of the coordinate vector fields $\partial_{r}$ with $r>k+\ell^{\prime}$ only. As in part (1), we can then modify these to fields $\eta_{k+\ell^{\prime}+1},\dots,\eta_{k+\ell}$ such that $\eta_{k+\ell^{\prime}+j}=\partial_{k+\ell^{\prime}+j}+\sum_{r>k+\ell}c_{rj}\partial_{r}$ for smooth functions $c_{rj}$.
@@
In particular, as in part (1) the vector fields $\eta_{k+\ell^{\prime}+1},\dots,\eta_{k+\ell}$ project to a local frame for $\mathcal{V}_{i^{\prime}}M$ that descends to a local frame for $T^{i^{\prime}}N/T^{i^{\prime}+1}N$. Now in the domain of our Frobenius chart, take a section of $\mathcal{V}_{i^{\prime}}M$ and a representative in $\Gamma\bigl{(}T^{i^{\prime}}_{P}M\bigr{)}$. Expanding this in terms of our local frame, we can drop all summands except the last $\ell-\ell^{\prime}$ ones without changing the projection to $\Gamma(\mathcal{V}_{i^{\prime}}M)$. Hence, we find a representative of the form $\eta=\sum_{j=1}^{\ell-\ell^{\prime}}c_{j}\eta_{k+\ell^{\prime}+j}$ and $[\partial_{i},\eta]=\sum_{j=1}^{\ell-\ell^{\prime}}(\partial_{i}\cdot c_{j})\eta_{k+\ell+j}$. The form of $\eta_{k+\ell+j}$ readily implies that this is a section of $T^{i^{\prime}+1}M$ if and only if it is zero and hence to $\partial_{i}\cdot c_{j}=0$ for any $i$ and $j$. But this is equivalent to each $c_{j}$ being independent of the first $k$ coordinates. Hence, $\eta$ descends to a local section of $T^{i^{\prime}}N/T^{i^{\prime}+1}N$ whose pullback visibly coincides with the original section of $\mathcal{V}_{i^{\prime}}M$. Conversely, in our frame, any such pullback has coordinate functions that descend to $N$. ∎
3.3 The distinguished class of relative tractor bundles
@@
In the setting of Section 2.3, let us assume that we have given a parabolic geometry ${(p\colon\mathcal{G}\!\to\!M,\omega)}$ whose torsion $\tau$ satisfies $\tau\bigl{(}T_{\rho}M,T^{i^{\prime}}_{P}M\bigr{)}\subset T^{i^{\prime}}_{P}M$ for all $i^{\prime}\leq 0$. Then $T_{\rho}M$ is involutive and by Theorem 3.4 for any local leaf space $\psi\colon U\to N$ for the induced foliation with connected fibers, we get a canonical filtration of the tangent bundle by the smooth subbundles $T^{i^{\prime}}N$. As usual, one can form the associated graded of this filtration, which has the form
@@
$\operatorname{gr}(TN)=\bigoplus_{i^{\prime}}\operatorname{gr}_{i^{\prime}}(TN)\qquad\text{with}\quad\operatorname{gr}_{i^{\prime}}(TN)=T^{i^{\prime}}N/T^{i^{\prime}+1}N\quad\text{and}\quad T^{0}N=N\times\{0\}.$
@@
Performing natural constructions with vector bundles, one always obtains induced filtrations of the resulting bundles, and forming the associated graded is compatible with the constructions. For example, the natural filtration of the cotangent bundle $T^{*}N$ has positive indices with the $j$th filtration component being the annihilator of $T^{-j+1}N$ so $TN=T^{1}N\supset T^{2}N\supset\cdots$. For the associated graded, this readily implies that $\operatorname{gr}_{j}(T^{*}N)$ is dual to $\operatorname{gr}_{-j}(TN)$. For filtered bundles $\bigl{(}E,E^{i}\bigr{)}$ and $\bigl{(}F,F^{j}\bigr{)}$, the components of the filtration on the tensor product $E\otimes F$ are defined by letting $(E\otimes F)^{\ell}$ be the span of the component $E^{i}\otimes F^{j}$ with $i+j=\ell$. For the associated graded, this means that $\operatorname{gr}_{\ell}(E\otimes F)=\bigoplus_{i+j=\ell}\operatorname{gr}_{i}(E)\otimes\operatorname{gr}_{j}(F)$. See [7, Section 3.1.1] for more details.
@@
Corollary 3.5 (to Theorems 3.3 and 3.4).
In the setting of Section 2.3, consider a parabolic geometry $(p\colon\mathcal{G}\to M,\omega)$ whose torsion $\tau$ satisfies $\tau\bigl{(}T_{\rho}M,T^{i^{\prime}}_{P}M\bigr{)}\subset T^{i^{\prime}}_{P}M$ for all $i^{\prime}\leq 0$. Let ${\psi\colon U\to N}$ be a local leaf space for the involutive distribution $T_{\rho}M=T^{0}_{P}M$ with connected fibers.
  • $(1)$
    For any $i^{\prime}$ we get $\mathcal{V}_{i^{\prime}}M|_{U}=\psi^{*}(\operatorname{gr}_{i^{\prime}}(TN))$. Applying any tensorial construction to the bundles $\mathcal{V}_{i^{\prime}}M$, one obtains a relative tractor bundle whose restriction to $U$ can be naturally identified with the pullback of the same construction applied to the bundles $\operatorname{gr}_{i^{\prime}}(TN)$.
  • $(2)$
    Suppose that the geometry even satisfies $\tau\bigl{(}T_{\rho}M,T^{i^{\prime}}_{P}M\bigr{)}\subset T^{i^{\prime}+1}_{P}M$ for all $i^{\prime}<0$. Then for any relative tractor bundle constructed as in $(1)$ a local section over $U$ is parallel for the relative tractor connection if and only if it is the pullback of a section of the corresponding bundle over $N$.
@@
(1) The first part immediately follows from Theorems 3.3 and 3.4. Applying a tensorial construction vector bundles associated to a principal bundle, one always obtains the bundle associated to the representation obtained from the corresponding construction on the inducing representations. Hence, any tensorial construction with the bundles $\mathcal{V}_{i^{\prime}}M$ leads to a relative tractor bundle. The last part then follows from the compatibility of tensorial constructions with the passage from filtered vector bundles to associated graded bundles as discussed above and with pullbacks.
@@
(2) Recall the construction of the relative tractor connection in terms of the fundamental derivative (or its relative analog) and the infinitesimal representation as described in Section 3.1. Naturality of the fundamental derivative (see [7, Section 1.5.8]) then implies that applying a tensorial construction to the bundles $\mathcal{V}_{i^{\prime}}M$, the relative tractor connection on the resulting relative tractor bundle coincides with the connection induced by the relative tractor connection on the bundles $\mathcal{V}_{i^{\prime}}M$. Knowing this, the general claim follows from Theorem 3.4. ∎
3.4 Example: Legendrean contact structures